# Think Twice before Adaptation: Improving Adaptability of DeepFake Detection via Online Test-Time Adaptation

### v1

@@ 5.1.2 Metrics
We use three evaluation metrics: accuracy (ACC), the area under the ROC curve (AUC), and average precision (AP). For each metric, higher values show better results. Notably, in the DF detection context, datasets inherently exhibit significant class imbalance with fake samples substantially dominating real ones [29], the AUC metric is more important as it remains robust to this problem.
@@ 5.1.3 Postprocessing Techniques
Following [6], we employ four postprocessing techniques: Gaussian blur, changes in color saturation, changes in color contrast, and resize: downsample the image by a factor then upsample it to the original resolution. At the inference time, test samples are applied to these operations with the intensity level increasing from $1$ to $5$. Details of postprocessing techniques and intensity levels are provided in Appendix C. Note that these postprocessing techniques are unknown to all models.
@@ 5.1.4 Baselines
For TTA, we compare our T2A method with SOTA methods, including TENT [23], MEMO [24], EATA [27], CoTTA [26], LAME [25], ViDA [28], and COME [7]. For DF detection, we employ the following DF detectors: EfficientNetB4 [46], F3Net [47], CORE [4], RECCE [11]. Details for these baselines are provided in Appendix C.
@@ 6 Conclusion
In this work, we introduce T2A, which improves the adaptability of DF detectors across two challenging scenarios: unknown postprocessing techniques and data distributions during inference time. Instead of solely relying on EM, T2A enables the model to explore alternative options before decision-making through NL with noisy pseudo-labels. We also provide a theoretical analysis to demonstrate that the proposed objective exhibits complementary behavior to EM. Through experiments, we show that T2A achieves higher adaptation performance compared to SoTA TTA approaches. Furthermore, when integrated with T2A, the resilience and generalization of DF detectors can be significantly improved without requiring additional training data or architectural modifications, making it particularly valuable for real-world deployments.
@@ Acknowledgments
Appendix for "Think Twice before Adaptation: Improving Adaptability of DeepFake Detection via Online Test-Time Adaptation"
@@ Appendix B More Experiments of Generation Artifacts
Figure 3 illustrates frequency spectra of the fake sample generated by StarGAN2. This fake sample is applied by 4 types of postprocessing operations, with the intensity level increasing from $1$ to $5$. Figure 2 shows the performance degradation of DF detectors under different types of postprocessing techniques across 5 intensity levels. Note that, in this experimental evaluation, both training and test samples are drawn from the same underlying data distribution (FaceForensics++ [39]), and only postprocessing operations are unseen during the testing phase.
@@ C.1 Datasets
We use FF++ [39] for training the source model (Xception) and other DF detectors. In this dataset, real videos collected from YouTube, which are then used to generate fake videos through four DF methods, including DeepFake, Face2Face, FaceSwap, and NeuralTexture. FF++ contains a total of 5000 videos, of which 1000 videos are sourced from YouTube.
C.2 Test datasets.
@@
  • CelebDF-v1 and CelebDF-v2 [40]: contain 998 real videos collected from 59 celebrities and 6434 fake videos improved by using techniques, such as higher resolution synthesis, color mismatch reduction, improved face mask, temporal flickering reduction. Videos in CelebDF datasets are variations in face sizes, orientations, lighting condiitons and backgrounds.
  • DeepFakeDetection (DFD) [41]: includes 363 real videos and 3000 fake videos.
  • DeepFake Detection Challenge Preview (DFDCP) [42]: consists of 1131 real videos of 66 individuals total and 4119 fake videos generated by multiple synthesis methods. Videos include varied lighting conditions, head poses, and backgrounds.
  • UADFV [43] which is composed of 98 real and fake videos from 49 different identities. This dataset mainly focuses on blinking, assisting in DF detection through physiological signals.
  • FaceShifter (FSh) [44]: includes a total of 2000 real and fake videos.
The image size of training and test samples are $256\times 256$ unless using Resize postprocessing (described in Sec. C.3). During the testing phase, individual frames extracted from these videos serve as our evaluation data. We use test sets of these datasets provided by [45].
(a) Resize
(b) Gaussian Blur
(c) Color Contrast
(d) Color Saturation
C.3 Intensity Levels of Postprocessing Techniques
In practice, both authentic and manipulated images frequently undergo various postprocessing operations. For real-world DF detection requirements, resilience to unknown postprocessing techniques is crucial. Following [6], we evaluate detector robustness across four fundamental postprocessing operations: Gaussian blur, resize, color saturation, and color contrast. For each operation, we implement five intensity levels based on standard corruption benchmarking practices [48]. Figure 4 shows an example of the five intensity levels of four types of postprocessing techniques.
@@ C.3 Intensity Levels of Postprocessing Techniques
C.4 Implementation Details
C.4.1 TTA baselines
For all TTA approaches, TENT [23], MEMO [24], EATA [27], CoTTA [26], LAME [25], ViDA [28], and COME [7], we follow all hyperparameters that are set in their Github unless it does not provide.
C.4.2 DF Detection baselines
Since the pre-trained models of EfficientNetB4 [46], F3Net [47], CORE [4], and RECCE [11] are not provided, we use that provided by [45].
C.4.3 Hyperparameters
@@
C.4.4 Analysis on Proposed Loss Functions
@@ D.4 Wall-clock running time of T2A
Figure 6: Average running time per iteration of TTA methods.

### v2

@@ 5.1.2 Metrics
We use three evaluation metrics: accuracy (ACC), the area under the ROC curve (AUC), and average precision (AP). For each metric, higher values show better results. Notably, in the DF detection context, datasets inherently exhibit significant class imbalance with fake samples substantially dominating real ones Layton et al. (2024), the AUC metric is more important as it remains robust to this problem.
@@ 5.1.3 Postprocessing Techniques
Following Chen et al. (2022), we employ four postprocessing techniques: Gaussian blur, changes in color saturation, changes in color contrast, and resize: downsample the image by a factor then upsample it to the original resolution. At the inference time, test samples are applied to these operations with the intensity level increasing from $1$ to $5$. Details of postprocessing techniques and intensity levels are provided in Appendix C. Note that these postprocessing techniques are unknown to all models.
@@ 5.1.4 Baselines
For TTA, we compare our T2A method with SOTA methods, including TENT Wang et al. (2020), MEMO Zhang et al. (2022), EATA Niu et al. (2022), CoTTA Wang et al. (2022), LAME Boudiaf et al. (2022), ViDA Liu et al. (2023a), and COME Zhang et al. (2024). For DF detection, we employ the following DF detectors: EfficientNetB4 Tan and Le (2019), F3Net Qian et al. (2020), CORE Ni et al. (2022), RECCE Cao et al. (2022). Details for these baselines are provided in Appendix C.
@@ 6 Conclusion
In this work, we introduce T2A, which improves the adaptability of DF detectors across two challenging scenarios: unknown postprocessing techniques and data distributions during inference time. Instead of solely relying on EM, T2A enables the model to explore alternative options before decision-making through NL with noisy pseudo-labels. We also provide a theoretical analysis to demonstrate that the proposed objective exhibits complementary behavior to EM. Through experiments, we show that T2A achieves higher adaptation performance compared to SoTA TTA approaches. Furthermore, when integrated with T2A, the resilience and generalization of DF detectors can be significantly improved without requiring additional training data or architectural modifications, making it particularly valuable for real-world deployments. However, since our method is based on backpropagation for updating parameters at inference time, it only works with end-to-end DF detectors that allow gradient flow throughout the model.
@@ Acknowledgments
Appendix for ”Think Twice before Adaptation: Improving Adaptability of DeepFake Detection via Online Test-Time Adaptation”
{tblr}
@@ Appendix B More Experiments of Generation Artifacts
Figure 3 illustrates frequency spectra of the fake sample generated by StarGAN2. This fake sample is applied by 4 types of postprocessing operations, with the intensity level increasing from $1$ to $5$. Figure 2 shows the performance degradation of DF detectors under different types of postprocessing techniques across 5 intensity levels. Note that, in this experimental evaluation, both training and test samples are drawn from the same underlying data distribution (FaceForensics++ Rossler et al. [2019]), and only postprocessing operations are unseen during the testing phase.
@@ C.1 Datasets
We use FF++ Rossler et al. [2019] for training the source model (Xception) and other DF detectors. In this dataset, real videos collected from YouTube, which are then used to generate fake videos through four DF methods, including DeepFake, Face2Face, FaceSwap, and NeuralTexture. FF++ contains a total of 5000 videos, of which 1000 videos are sourced from YouTube.
C.1.2 Test datasets.
@@
  • CelebDF-v1 and CelebDF-v2 Li et al. [2020b]: contain 998 real videos collected from 59 celebrities and 6434 fake videos improved by using techniques, such as higher resolution synthesis, color mismatch reduction, improved face mask, temporal flickering reduction. Videos in CelebDF datasets are variations in face sizes, orientations, lighting condiitons and backgrounds.
  • DeepFakeDetection (DFD) Google [2019]: includes 363 real videos and 3000 fake videos.
  • DeepFake Detection Challenge Preview (DFDCP) Dolhansky [2019]: consists of 1131 real videos of 66 individuals total and 4119 fake videos generated by multiple synthesis methods. Videos include varied lighting conditions, head poses, and backgrounds.
  • UADFV Li et al. [2018] which is composed of 98 real and fake videos from 49 different identities. This dataset mainly focuses on blinking, assisting in DF detection through physiological signals.
  • FaceShifter (FSh) Li et al. [2020a]: includes a total of 2000 real and fake videos.
The image size of training and test samples are $256\times 256$ unless using Resize postprocessing (described in Sec. C.2). During the testing phase, individual frames extracted from these videos serve as our evaluation data. We use test sets of these datasets provided by Yan et al. [2023].
C.2 Intensity Levels of Postprocessing Techniques
In practice, both authentic and manipulated images frequently undergo various postprocessing operations. For real-world DF detection requirements, resilience to unknown postprocessing techniques is crucial. Following Chen et al. [2022], we evaluate detector robustness across four fundamental postprocessing operations: Gaussian blur, resize, color saturation, and color contrast. For each operation, we implement five intensity levels based on standard corruption benchmarking practices Hendrycks and Dietterich [2019]. Figure 6 shows an example of the five intensity levels of four types of postprocessing techniques.
@@ C.2 Intensity Levels of Postprocessing Techniques
C.3 Implementation Details
C.3.1 TTA baselines
For all TTA approaches, TENT Wang et al. [2020], MEMO Zhang et al. [2022], EATA Niu et al. [2022], CoTTA Wang et al. [2022], LAME Boudiaf et al. [2022], ViDA Liu et al. [2023a], and COME Zhang et al. [2024], we follow all hyperparameters that are set in their Github unless it does not provide.
C.3.2 DF Detection baselines
Since the pre-trained models of EfficientNetB4 Tan and Le [2019], F3Net Qian et al. [2020], CORE Ni et al. [2022], and RECCE Cao et al. [2022] are not provided, we use that provided by Yan et al. [2023].
C.3.3 Hyperparameters
@@
Figure 5: Average running time per iteration of TTA methods.
@@ D.1.2 Analysis on Proposed Loss Functions
(a) Resize
(b) Gaussian Blur
(c) Color Contrast
(d) Color Saturation